# Power sector effects of alternative options for de-fossilizing heavy-duty vehicles: go electric, and charge smartly

### v3

@@ Abstract
In the passenger car segment, battery-electric vehicles (BEV) have emerged as the most promising option to de-fossilize transportation. For heavy-duty vehicles (HDV), the technology space still appears to be more open. Aside from BEV, electric road systems (ERS) for dynamic power transfer are discussed, as well as indirect electrification with trucks that use hydrogen fuel cells or e-fuels. Here we investigate the power sector implications of these alternative options. We apply an open-source capacity expansion model to future scenarios of Germany with high renewable energy shares, drawing on detailed route-based truck traffic data. Results show that power sector costs are lowest for flexibly charged BEV that also carry out vehicle-to-grid operations, and highest for HDV using e-fuels. If BEV and ERS-BEV are not charged in an optimized way, power sector costs increase, but are still substantially lower than in scenarios with hydrogen or e-fuels. This is a consequence of the poor energy efficiency of indirect HDV electrification, which outweighs its temporal flexibility benefits. We further find a higher use of solar photovoltaic energy for BEV and ERS-BEV, while hydrogen and e-fuel HDV lead to a higher use of wind power and fossil electricity generation. Results are qualitatively robust in sensitivity analyses without the European interconnection or with capacity limits for wind power expansion.
@@
Heavy-duty vehicles, battery-electric vehicles, catenary, hydrogen, power sector modeling
@@ 1 Introduction
Making energy consumption climate neutral in all end-use sectors is of paramount importance for mitigating climate change de_coninck_strengthening_2018. A key strategy for achieving this is to substitute fossil fuels by renewable electricity, facilitated by direct or indirect electrification of end uses in mobility, heating, and industrial applications shukla_contribution_2022. In the transportation sector, battery-electric vehicles (BEV) have emerged as the most promising option for de-fossilizing the passenger car segment. Already today, BEV can lead to sizeable greenhouse gas emission reductions compared to internal combustion engines hoekstra_underestimated_2019, which will further increase when the electricity mix becomes cleaner. In many countries, markets for electric passenger cars have been soaring in the past years, and are expected to continue to grow strongly in the near future iea_global_2022. For heavy-duty vehicles (HDV), however, the technology space still appears to be more open. While the feasibility of pure battery-electric HDV has been assessed to be increasing nykvist_feasibility_2021, they compete with other options. This includes electric road systems (ERS), which allow for dynamic power transfer to electric vehicles on the road (boltze_insights_2020; speth_comparing_2021); trucks with hydrogen fuel cells; or conventional HDV with internal combustion engines that use power-to-liquid (PtL), also referred to as e-fuels, which are produced mostly with renewable electricity hannula_nearterm_2019; lajevardi_simulating_2022; plotz_hydrogen_2022; li_transition_2022.
@@
These options of direct or indirect electrification of HDV have different properties concerning, on the one hand, energy efficiency, and, on the other hand, temporal flexibility of electricity use. For example, direct electrification via BEV is more energy efficient compared to indirect electrification via electrolysis-based hydrogen or e-fuels ueckerdt_2021; lajevardi_simulating_2022. Yet, the temporal flexibility of BEV may be constrained by charging availability and limited battery capacities, as vehicle batteries are costly and heavy. In contrast, indirect electrification via hydrogen or e-fuels may entail large-scale and low-cost storage options taljegard_2017; stoeckl_2021, but the overall energy efficiency of these supply chains is lower compared to BEV. Temporal power sector flexibility becomes increasingly important with growing shares of renewables, as the potential for firm renewable generation such as hydropower, bioenergy, or geothermal power is limited in many countries. In contrast, wind and solar power potentials are often abundant, but they have variable generation profiles that depend on weather conditions and daily and seasonal cycles (lopezprol_economics_2021). Integrating growing shares of variable renewable energy sources thus requires an increasing use of temporal and spatial flexibility options in the power sector. Flexibility may be provided on the supply side (e.g., dispatchable generators) or the demand side (e.g., flexible loads), and via energy storage or transmission (kondziella_flexibility_2016; schill_joule_2020).
@@
Against this background, we investigate the power sector implications of different options for (in-)directly electrifying HDV, particularly focusing on the trade-off between energy efficiency and temporal flexibility. To do so, we apply an open-source capacity expansion model zerrahn_long-run_2017; gaete_2021 to scenarios of the Central European power sector for the year 2030 with high renewable energy shares. We focus on the domestic traffic of HDV in Germany with a gross vehicle weight above 26 tonnes, drawing on a detailed data set of truck trips on inner-German routes. We include stationary-charged BEV trucks as well as hybrid battery-catenary trucks as a particular example of an electric road system technology (ERS-BEV) [^1] In general, there are three options available for ERS: Overhead catenary lines, conductor rails, and inductive power transfer. We have opted for the first option as it is the technology currently under discussion in Germany with three test tracks in place (Hacker2023)., fuel-cell hydrogen electric trucks (FCEV), and such with internal diesel combustion engines powered by e-fuels (ICEV PtL), see Figure 1. For hydrogen and e-fuels, we assume domestic supply chains. We compare the power sector costs [^2] Power sector costs comprise fixed and variable costs of all electricity generation and storage technologies over a full year. They do not include the costs of HDV electrification infrastructure, e.g., overhead lines, charging stations, or filling stations for hydrogen or PtL. of these options, as well as their repercussions on the optimal power plant fleet and its use.
@@
While there is a broad literature on the potential power sector impacts of battery-electric passenger cars richardson_electric_2013; muratori_shape_2020; mangipinto_impact_2022, according research for electric HDV is sparse. A common finding of analyses focusing on passenger cars is that controlled charging can help to avoid problematic load peaks, to integrate renewable electricity generation, and to lower overall system costs schill_power_2015; gnann_2018; sadeghian_a_2022. In particular, feeding back electricity from car batteries to the system (vehicle-to-grid, V2G) can help to balance the diurnal variability of solar photovoltaic (PV) generation brown_2018.
How electrified HDV interact with the power sector is hardly explored so far. Some analyses investigate grid or peak load impacts without using power sector models. For example, an analysis of the distribution grid impacts of BEV-HDV with depot charging finds that many existing substations in the US would allow for massive HDV charging without grid upgrades borlaug_heavyduty_2021. Earlier studies conclude that local grid impacts of electrified HDV near logistics centres and along major roads may be substantial in Switzerland and Finland (liimatainen_2019), or that ERS-HDV could strongly increase the peak load in Norway taljegard_2017. Using capacity expansion power sector models, other analyses come to deviating conclusions on the impacts of electric HDV on the overall load profile and the wider power sector; in a Scandinavian-German case study that includes ERS-HDV in a subset of scenarios, effects are pronounced taljegard_impacts_2019, while in a European model study the impacts are moderate plotz_impact_2019. Importantly, both analyses assume exogenous, i.e., inflexible, charging profiles of HDV and do not temporally optimize their (dis-)charging operations. Given this research gap, we deliberately focus our analysis on the power sector effects of electrified HDV, modelling in detail their repercussions on optimal capacities and their hourly use. In doing so, we complement other strand of the literature which have focused on the total costs of ownership (TCO) nykvist_feasibility_2021; hunter_spatial_2021; samet_2024, energy efficiencynoll_analyzing_2022, or potential health and climate impacts of electric HDV lin_2021.
@@ 2 Results
In the main part of this paper, we show results for a setting where Germany is interconnected with its neighbors, defined as the baseline model specification. Each country is modeled as a single node. While this configuration can be considered to be a policy-relevant scenario for 2030, where electric HDV in Germany may benefit from the flexibility provided by the European interconnection, we also provide results for a sensitivity analysis where the German power sector is modeled in isolation in the Supplemental Information (SI.4.1). [^3] In the SI, we also show further results for baseline assumptions (section SI.3), a sensitivity analysis that explores the effects of an upper bound for wind power capacity (section SI.4.2), and another sensitivity with alternative charging availability assumptions (section SI.4.3).
@@ 2.1 Lowest power sector costs and electricity prices for BEV with V2G
Compared to the reference case without electrified trucks, yearly power sector costs increase in all scenarios with electrified HDV (Figure 2, upper panel). The cost of the additional electricity demand induced by HDV thus always outweighs their potential flexibility benefits. Cost effects, however, vary considerably between different options. Flexible BEV with V2G incur the lowest additional power sector costs of 1.8 bn Euros/year (around 5,600 Euros/year per vehicle), followed by optimally charged BEV without V2G with 2.3 bn Euros/year (around 7,200 Euros/year per vehicle). If BEV charging is not optimized, system costs are markedly higher, with 3.8 bn Euros/year (11,900 Euros/year per vehicle). Results are qualitatively similar for ERS-BEV, but on a slightly higher cost level. The differences between the three ERS-BEV cases are less pronounced than for pure BEV, as their temporal flexibility potential is much smaller. The battery capacity of an ERS-BEV is only around a quarter of that of a pure BEV (655 kWh usable capacity per truck vs. 181 kWh), so ERS-BEV can make less use of hours with low electricity prices (compare also Figure SI.9). In contrast, power sector costs are higher for FCEV (12.6 or 12.7 bn Euros/year, for decentralized or centralized hydrogen provision, i.e. around 39,700 Euros/year per vehicle) and even more so for PtL (16.8 bn Euros/year, or 52,700 Euros/year per vehicle). This is a consequence of high conversion losses of hydrogen and PtL supply chains and vehicles drive trains. Because of these losses, the two hydrogen supply chains increase the electricity demand more than twice as much as the battery-electric options. The electricity demand of PtL-HDV is nearly four times as high as in the case of BEV. Notably, the cost differences between BEV and ERS-BEV scenarios are much smaller than the differences between these direct-electric options and indirect electrification via hydrogen or PtL.
@@
We also evaluate average yearly wholesale electricity prices for HDV electricity (Figure 2, lower panel). [^4] We interpret the dual variables of the model’s energy balance as hourly wholesale prices, which is common practice in energy modeling (brown_decreasing_2021; lopezprol_economics_2021). This allows to separate the differences in overall electricity consumption of the various HDV options from their ability to make use of low-cost electricity. Average prices are calculated by multiplying hourly wholesale prices of electricity consumed by the different HDV options or fed back to the grid with respective hourly quantities, summing up over the whole year, and dividing by the overall electricity consumption of the fleet. That is, the numbers also account for revenues of electricity sold via V2G, assuming that HDV operators receive the respective hourly wholesale price whenever they feed electricity back to the grid. BEV with V2G accordingly face the lowest average electricity prices, as these also benefit from revenues of feeding back to the grid, followed by BEV without V2G.
@@ 2.2 Capacity and dispatch effects
The upper panel of Figure 3 shows optimal generation capacities in the reference (on the left) and the changes induced by HDV (on the right) in Germany. In the reference, variable solar PV and onshore wind power dominate the capacity mix. [^5] In the Supplemental Information, we discuss the effects of including an upper bound for wind power investments by the year 2030, see section SI.4. These are complemented by smaller firm capacities of natural gas and bioenergy. Optimal capacity additions related to the electrification of HDV are predominantly a mix of solar PV and onshore wind power, as these are the lowest-cost options available. Flexible BEV and ERS-BEV lead to the highest PV shares in the capacity additions, especially if combined with V2G. BEV trucks with V2G serve as a short-duration electricity storage option with a low energy-to-power ratio and many yearly cycles, which favors the expansion of solar PV. The potential for solar PV integration via V2G is less pronounced for ERS-BEV than for BEV, because ERS-BEV have smaller batteries.
[…]
@@
Yearly electricity generation in the reference, as well as the changes induced by different HDV scenarios, are shown in the lower panel of Figure 3. The picture is generally similar to optimal capacity changes, but the share of wind power in additional electricity generation tends to be higher because of higher full-load hours as compared to solar PV. Flexible BEV, particularly if combined with V2G, cause the highest increase in PV generation, which can increase the renewable share of the German power sector to nearly 84% (section SI.3.3). Centralized hydrogen or PtL supply chains, which require much more electricity than BEV and ERS-BEV, use their temporal flexibility to also increase electricity imports and the use of the most energy-efficient natural gas-fired power plants (CCGT). Note that this flexibility comes from the option of low-cost hydrogen or PtL storage. This is different in the case of temporally inflexible hydrogen production at filling stations (FCEV On-Site). As imports and CCGT are less available in the hours of decentralized hydrogen production, more wind and solar generation in combination with long-duration electricity storage are used, which also increases the share of renewables in the overall system to nearly 83%. Higher utilization of natural gas also translates into increasing direct carbon emissions (section SI.3.4). The effects of BEV and ERS-BEV on CO2 emissions are generally lower, and may even become negative when combined with V2G. This is because V2G allows for an increase in solar photovoltaic capacity and a higher renewable energy share, which can slightly reduce overall carbon emissions compared to the reference case without electrified HDV.

### v4

# Power sector effects of alternative options for de-fossilizing heavy-duty vehicles – go electric, and charge smartly

@@ Abstract
Various options are discussed to de-fossilize heavy-duty vehicles (HDV), including battery-electric vehicles (BEV), electric road systems (ERS), and indirect electrification via hydrogen fuel cells or e-fuels. We investigate their power sector implications in future scenarios of Germany with high renewable energy shares, using an open-source capacity expansion model and route-based truck traffic data. Power sector costs are lowest for flexibly charged BEV that also carry out vehicle-to-grid operations, and highest for e-fuels. If BEV and ERS-BEV are not optimally charged, power sector costs increase, but are still substantially lower than in scenarios with hydrogen or e-fuels. This is because indirect electrification is less energy efficient, which outweighs potential flexibility benefits. BEV and ERS-BEV favor solar photovoltaic energy, while hydrogen and e-fuels favor wind power and increase fossil electricity generation. Results remain qualitatively robust in sensitivity analyses.
@@
Electric trucks , battery-electric vehicles , catenary , hydrogen , power sector modeling
journal: arXiv
[…]
linewidth=0pt, innerleftmargin=0pt, innerrightmargin=0pt, innertopmargin=0pt, innerbottommargin=0pt
@@ 1 Introduction
Making energy consumption climate neutral in all end-use sectors is of paramount importance for mitigating climate change [de_coninck_strengthening_2018]. A key strategy for achieving this is to substitute fossil fuels by renewable electricity, facilitated by direct or indirect electrification of end uses in mobility, heating, and industrial applications [shukla_contribution_2022]. In the transportation sector, battery-electric vehicles (BEV) have emerged as the most promising option for de-fossilizing the passenger car segment. Already today, BEV can lead to sizeable greenhouse gas emission reductions compared to internal combustion engines [hoekstra_underestimated_2019], which will further increase when the electricity mix becomes cleaner. In many countries, markets for electric passenger cars have been soaring in the past years, and are expected to continue to grow strongly in the near future [iea_global_2023]. For heavy-duty vehicles (HDV), however, the technology space still appears to be more open. While the feasibility of pure battery-electric HDV has been assessed to be increasing [nykvist_feasibility_2021], they compete with other options. This includes electric road systems (ERS), which allow for dynamic power transfer to electric vehicles on the road [boltze_insights_2020, speth_comparing_2021]; trucks with hydrogen fuel cells; or conventional HDV with internal combustion engines that use power-to-liquid (PtL), also referred to as e-fuels, which are produced mostly with renewable electricity [hannula_nearterm_2019, lajevardi_simulating_2022, plotz_hydrogen_2022, li_transition_2022].
@@
These options of direct or indirect electrification of HDV have different properties concerning, on the one hand, energy efficiency, and, on the other hand, temporal flexibility of electricity use. For example, direct electrification via BEV is more energy efficient compared to indirect electrification via electrolysis-based hydrogen or e-fuels [ueckerdt_2021, lajevardi_simulating_2022]. Yet, the temporal flexibility of BEV may be constrained by charging availability and limited battery capacities, as vehicle batteries are costly and heavy. In contrast, indirect electrification via hydrogen or e-fuels may entail large-scale and low-cost storage options [taljegard_2017, stoeckl_2021], but the overall energy efficiency of these supply chains is lower compared to BEV. Temporal power sector flexibility becomes increasingly important with growing shares of renewables, as the potential for firm renewable generation such as hydropower, bioenergy, or geothermal power is limited in many countries. In contrast, wind and solar power potentials are often abundant, but they have variable generation profiles that depend on weather conditions and daily and seasonal cycles [lopezprol_economics_2021]. Integrating growing shares of variable renewable energy sources thus requires an increasing use of temporal and spatial flexibility options in the power sector. Flexibility may be provided on the supply side (e.g., dispatchable generators) or the demand side (e.g., flexible loads), and via energy storage or transmission [kondziella_flexibility_2016, schill_joule_2020].
@@
Against this background, we investigate the power sector implications of different options for (in-)directly electrifying HDV, particularly focusing on the trade-off between energy efficiency and temporal flexibility. To do so, we apply an open-source capacity expansion model [zerrahn_long-run_2017, gaete_2021] to scenarios of the Central European power sector for the year 2030 with high renewable energy shares. We focus on the domestic traffic of HDV in Germany with a gross vehicle weight above 26 tonnes, drawing on a detailed data set of truck trips on inner-German routes. We include stationary-charged BEV trucks as well as hybrid battery-catenary trucks as a particular example of an electric road system technology (ERS-BEV) [^1] In general, there are three options available for ERS: Overhead catenary lines, conductor rails, and inductive power transfer. We have opted for the first option as it is the technology currently under discussion in Germany with three test tracks in place [Hacker2023]., fuel-cell hydrogen electric trucks (FCEV), and such with internal diesel combustion engines powered by e-fuels (ICEV PtL), see Figure 1. For hydrogen and e-fuels, we assume domestic supply chains. We compare the power sector costs [^2] Power sector costs comprise fixed and variable costs of all electricity generation and storage technologies over a full year. They do not include the costs of HDV electrification infrastructure, e.g., overhead lines, charging stations, or filling stations for hydrogen or PtL. of these options, as well as their repercussions on the optimal power plant fleet and its use.
{mdframed}
@@
While there is a broad literature on the potential power sector impacts of battery-electric passenger cars [richardson_electric_2013, muratori_shape_2020, mangipinto_impact_2022], according research for electric HDV is sparse. A common finding of analyses focusing on passenger cars is that controlled charging can help to avoid problematic load peaks, to integrate renewable electricity generation, and to lower overall system costs [schill_power_2015, gnann_2018, sadeghian_a_2022]. In particular, feeding back electricity from car batteries to the system (vehicle-to-grid, V2G) can help to balance the diurnal variability of solar photovoltaic (PV) generation [brown_2018].
How electrified HDV interact with the power sector is hardly explored so far. Some analyses investigate grid or peak load impacts without using power sector models. For example, an analysis of the distribution grid impacts of BEV-HDV with depot charging finds that many existing substations in the US would allow for massive HDV charging without grid upgrades [borlaug_heavyduty_2021]. Earlier studies conclude that local grid impacts of electrified HDV near logistics centres and along major roads may be substantial in Switzerland and Finland [liimatainen_2019], or that ERS-HDV could strongly increase the peak load in Norway [taljegard_2017]. Using capacity expansion power sector models, other analyses come to deviating conclusions on the impacts of electric HDV on the overall load profile and the wider power sector; in a Scandinavian-German case study that includes ERS-HDV in a subset of scenarios, effects are pronounced [taljegard_impacts_2019], while in a European model study the impacts are moderate [plotz_impact_2019]. Importantly, both analyses assume exogenous, i.e., inflexible, charging profiles of HDV and do not temporally optimize their (dis-)charging operations. Given this research gap, we deliberately focus our analysis on the power sector effects of electrified HDV, modelling in detail their repercussions on optimal capacities and their hourly use. In doing so, we complement other strand of the literature which have focused on the total costs of ownership (TCO) [nykvist_feasibility_2021, hunter_spatial_2021, samet_2024], energy efficiency[noll_analyzing_2022], or potential health and climate impacts of electric HDV [lin_2021].
@@ 2 Results
In the main part of this paper, we show results for a setting where Germany is interconnected with its neighbors, defined as the baseline model specification. Each country is modeled as a single node. While this configuration can be considered to be a policy-relevant scenario for 2030, where electric HDV in Germany may benefit from the flexibility provided by the European interconnection, we also provide results for a sensitivity analysis where the German power sector is modeled in isolation in the Supplemental Information (S.2.6). [^3] In the SI, we also show further results for baseline assumptions (sections S.2.1-S.2.5), a sensitivity analysis that explores the effects of an upper bound for wind power capacity (section S.2.7), and another sensitivity with alternative charging availability assumptions (section S.2.8).
@@ 2.1 Lowest power sector costs and electricity prices for BEV with V2G
Compared to the reference case without electrified trucks, yearly power sector costs increase in all scenarios with electrified HDV (Figure 2, upper panel). The cost of the additional electricity demand induced by HDV thus always outweighs their potential flexibility benefits. Cost effects, however, vary considerably between different options. Flexible BEV with V2G incur the lowest additional power sector costs of 1.8 bn Euros/year (around 5,600 Euros/year per vehicle), followed by optimally charged BEV without V2G with 2.3 bn Euros/year (around 7,200 Euros/year per vehicle). If BEV charging is not optimized, system costs are markedly higher, with 3.8 bn Euros/year (11,900 Euros/year per vehicle). Results are qualitatively similar for ERS-BEV, but on a slightly higher cost level. The differences between the three ERS-BEV cases are less pronounced than for pure BEV, as their temporal flexibility potential is much smaller. The battery capacity of an ERS-BEV is only around a quarter of that of a pure BEV (655 kWh usable capacity per truck vs. 181 kWh), so ERS-BEV can make less use of hours with low electricity prices (compare also Figure S.11). In contrast, power sector costs are higher for FCEV (12.6 or 12.7 bn Euros/year, for decentralized or centralized hydrogen provision, i.e. around 39,700 Euros/year per vehicle) and even more so for PtL (16.8 bn Euros/year, or 52,700 Euros/year per vehicle). This is a consequence of high conversion losses of hydrogen and PtL supply chains and vehicles drive trains. Because of these losses, the two hydrogen supply chains increase the electricity demand more than twice as much as the battery-electric options. The electricity demand of PtL-HDV is nearly four times as high as in the case of BEV. Notably, the cost differences between BEV and ERS-BEV scenarios are much smaller than the differences between these direct-electric options and indirect electrification via hydrogen or PtL.
@@
We also evaluate average yearly wholesale electricity prices for HDV electricity (Figure 2, lower panel). [^4] We interpret the dual variables of the model’s energy balance as hourly wholesale prices, which is common practice in energy modeling [brown_decreasing_2021, lopezprol_economics_2021]. This allows to separate the differences in overall electricity consumption of the various HDV options from their ability to make use of low-cost electricity. Average prices are calculated by multiplying hourly wholesale prices of electricity consumed by the different HDV options or fed back to the grid with respective hourly quantities, summing up over the whole year, and dividing by the overall electricity consumption of the fleet. That is, the numbers also account for revenues of electricity sold via V2G, assuming that HDV operators receive the respective hourly wholesale price whenever they feed electricity back to the grid. BEV with V2G accordingly face the lowest average electricity prices, as these also benefit from revenues of feeding back to the grid, followed by BEV without V2G.
@@ 2.2 Capacity and dispatch effects
The upper panel of Figure 3 shows optimal generation capacities in the reference (on the left) and the changes induced by HDV (on the right) in Germany. In the reference, variable solar PV and onshore wind power dominate the capacity mix. [^5] In the Supplemental Information, we discuss the effects of including an upper bound for wind power investments by the year 2030, see section S.2.7. These are complemented by smaller firm capacities of natural gas and bioenergy. Optimal capacity additions related to the electrification of HDV are predominantly a mix of solar PV and onshore wind power, as these are the lowest-cost options available. Flexible BEV and ERS-BEV lead to the highest PV shares in the capacity additions, especially if combined with V2G. BEV trucks with V2G serve as a short-duration electricity storage option with a low energy-to-power ratio and many yearly cycles, which favors the expansion of solar PV. The potential for solar PV integration via V2G is less pronounced for ERS-BEV than for BEV, because ERS-BEV have smaller batteries.
HDV options that are temporally less flexible or that require more electricity favor higher onshore wind power capacities. In this case, the wind power generation profile fits better to HDV electricity demand than solar PV. Offshore wind power is not added here because of unfavorable cost assumptions compared to onshore wind power. [^6] Given new policy targets and the outcomes of current tenders for offshore wind power capacity in Germany, offshore wind power may play a larger role than assumed here. FCEV and PtL options have the highest capacity needs because of substantial conversion losses. On-site electrolysis further requires an addition of long-duration electricity storage capacity (power-to-gas-to-power) of 9.4 GW and 6.0 TWh (see section S.2.1, Figure S.6). This is necessary to compensate for the temporal inflexibility of on-site electrolyzers (FigureS.7) without large-scale hydrogen storage, which are forced to operate with a similar time profile as hydrogen demand. Among inflexibly charged battery-electric HDV options, BEV require larger investments in stationary lithium-ion batteries than ERS-BEV. Largely due to a CO2 price of 100 Euros/ton, hard coal, lignite or oil generators are not used in any of the scenarios, while natural gas (CCGT) capacity is always at the assumed upper limit of 17.6 GW (compare section 4). Up to 5 GW of open cycle gas turbines (OCGT) are added in some scenarios, especially for HDV options with low temporal flexibility (BEV Inflex, ERS-BEV Inflex and FCEV On-Site).
@@
Yearly electricity generation in the reference, as well as the changes induced by different HDV scenarios, are shown in the lower panel of Figure 3. The picture is generally similar to optimal capacity changes, but the share of wind power in additional electricity generation tends to be higher because of higher full-load hours as compared to solar PV. Flexible BEV, particularly if combined with V2G, cause the highest increase in PV generation, which can increase the renewable share of the German power sector to nearly 84% (section S.2.3, Figure S.8). Centralized hydrogen or PtL supply chains, which require much more electricity than BEV and ERS-BEV, use their temporal flexibility to also increase electricity imports and the use of the most energy-efficient natural gas-fired power plants (CCGT). Note that this flexibility comes from the option of low-cost hydrogen or PtL storage. This is different in the case of temporally inflexible hydrogen production at filling stations (FCEV On-Site). As imports and CCGT are less available in the hours of decentralized hydrogen production, more wind and solar generation in combination with long-duration electricity storage are used, which also increases the share of renewables in the overall system to nearly 83%. Higher utilization of natural gas also translates into increasing direct carbon emissions (section S.2.4, Figure S.9). The effects of BEV and ERS-BEV on CO2 emissions are generally lower, and may even become negative when combined with V2G. This is because V2G allows for an increase in solar photovoltaic capacity and a higher renewable energy share, which can slightly reduce overall carbon emissions compared to the reference case without electrified HDV.